# Magnon transmission across $\nu=1|-1|1$ mono-layer graphene junction as a probe of electronic structure

### v1

@@ V.3 Magnon transmission at very small $E_{V}$
First, we consider the case where $u_{z}^{0}>|u_{xy}^{0}|$. We illustrate this case with the parameter choices $u_{z}^{0}=4,~{}u_{\perp}^{0}=-3,~{}E_{Z}=0.5,~{}E_{V}=0.1$. The HF one-body energy in the $\nu=1$ bulk are shown in Fig. 15. As can be seen, the ordering $E_{K\uparrow}>E_{K\downarrow}>E_{K^{\prime}\uparrow}$ is preserved for all $B_{\perp}$.
@@ VI Conclusions, caveats, and open questions
In this work, we study the transmission of spin magnons across a graphene $1|-1|1$ system. Our main motivation for studying this particular setup is to obtain knowledge about the ratio of lattice-scale, ultra-short-range anisotropic couplings $u_{z}$ and $u_{\perp}$. In physical graphene samples it is believed that $u_{z}>0$ while $u_{\perp}<0$, and that the two have roughly the same magnitude. Furthermore, their ratio can be altered by Landau-level mixing [67, 68, 69, 70] and the screening environment [71]. Pinning down this ratio and its evolution with screening and other tuning parameters would be invaluable in determining the phases of the $\nu=0$ system, which is yet to be fully understood.
In the setup, we used the bulk ground states for $\nu=\pm 1$. In contrast to $\nu=0$, where the nature of the phase is not fully understood, the ground state at $\nu=\pm 1$ in the physical range of couplings with both Zeeman and sublattice potential present is known to be a valley-polarized ferromagnet. This makes it easier to determine the values of the couplings themselves. At perpendicular field, $E_{Z}$ is known, and in principle, one can determine $E_{V}$ by measuring the gap at charge neutrality at $B_{\perp}=0$. A great advantage of this setup is that the in situ tuning parameter $B_{\perp}$ allows us to alter the ratio of $E_{V}$ with respect to the other couplings.
@@
The second important fact in understanding our results is that in the physical region of parameters, $u_{z}>0,~{}u_{\perp}<0$, when $u_{z}>|u_{\perp}|$ the system prefers to break the spin $U(1)$ symmetry at the interfaces, while in the opposite case $u_{z}<|u_{\perp}|$ the system prefers to break the valley $U(1)$ symmetry (for some particular ${B_{\perp}}$). Since the interfaces between $\nu=1$ and $\nu=-1$ can very roughly be thought of as miniature regions of $\nu=0$, this is consistent with the fact that in the corresponding regions the $\nu=0$ bulk ground state breaks exactly those symmetries.
@@
Keeping these two facts in mind, we can easily understand the cases that we considered in Section V with $u_{z}>|u_{\perp}|$. There is no valley rotation in these cases, and thus the magnon transmission drops to zero at high energy. The more interesting case is $u_{z}<|u_{\perp}|$, believed to occur for unscreened or lightly screened samples [71]. Here a crucial role is played by the valley Zeeman field $E_{V}$. For vanishing $E_{V}\to 0$ the physics of magnon transmission was analyzed by Wei et al[47], and there is nearly perfect transmission at high energy. Focusing on $E_{V}$ moderate to large, at small $B_{\perp}$, $E_{V}>|u_{\alpha}|,E_{z}$. In this case, the interfaces do not break any symmetries, and the magnon transmission vanishes at high energies. However, there is a threshold $B_{\perp}$ at which the couplings $u_{\alpha}$ become dominant over $E_{V}$, beyond which the interfaces break the valley $U(1)$ symmetry, restoring nearly perfect transmission at high energies. This threshold $B_{\perp}$ depends not only on the couplings, but also on the width of the intermediate $\nu=-1$ region, and can be estimated in our model given the sample geometry. This is one of the main results of our work, because this threshold field provides quantitative information about the coupling constants.
@@
Let us turn to some of the assumptions that underlie our approach. We have assumed ultra-short-range interactions throughout. Relaxing this assumption in the physical region of parameters $u_{z}>0,~{}u_{\perp}<0$ does not change the phases of $\nu=\pm 1$. However, introducing interactions beyond ultra-short-range does produce new phases at $\nu=0$. Because of our sharp interfaces, the $\nu=0$ regions here are fairly narrow, and we believe interactions beyond USR will not have any qualitative effect on our results. Secondly, we have focused on incident magnons with $q_{y}=0$. In Appendix A, we show some results for incident magnons with $q_{y}\neq 0$, which look qualitatively similar to our results in Section V. Thirdly, we have ignored disorder and finite temperature effects. Disorder can induce the magnons to scatter elastically, thereby reducing the transmission. At $T\neq 0$, thermally generated collective modes will be present in the system, and could scatter the electrically generated magnons inelastically.
@@
There are many open questions that could in principle be addressed by a detailed analysis of magnon transmission. The most important is the $\nu=0$ state, which remains to be completely understood. It is believed that Landau level mixing leads to the interactions acquiring a range of the magnetic length $\ell$ [41, 42, 71]. Introducing such interactions leads to the appearance of new phases which manifest the coexistence of CAF and bond order, and are separated from the bond-ordered and CAF states by second-order phase transitions [41, 42, 44]. It should be possible to vary the screening to make the system traverse this coexistence phase. Presumably, the magnon transmission properties of this phase differ from that of the standard CAF phase. Furthermore, fractional quantum Hall phases near $\nu=0$ also display a rich variety of phases in the physical regime of parameters, which could be explored via magnon transmission [29]. We hope to address these and other such questions in the near future.
@@ Acknowledgements.
The authors would like to thank Nemin Wei and Chunli Huang for illuminating discussions. SJD would like to acknowledge both Harish-Chandra Research Institute, Prayagraj, and the cluster computing facilities at the International Center for Theoretical Sciences (ICTS), Bangalore. He also acknowledges ICTS for hospitality and visitor funding. SR and GM would like to thank the VAJRA scheme of SERB, India for its support. GM thanks the Aspen Center for Physics (NSF Grant No. 1066293) for its hospitality.

### v2

@@ V.3 Magnon transmission at very small $E_{V}$
First, we consider the case where $u_{z}^{0}>|u_{xy}^{0}|$. We illustrate this case with the parameter choices $u_{z}^{0}=4,~{}u_{xy}^{0}=-3,~{}E_{Z}=0.5,~{}E_{V}=0.1$. The HF one-body energy in the $\nu=1$ bulk are shown in Fig. 15. As can be seen, the ordering $E_{K\uparrow}>E_{K\downarrow}>E_{K^{\prime}\uparrow}$ is preserved for all $B_{\perp}$.
@@ VI Conclusions, caveats, and open questions
In this work, we study the transmission of spin magnons across a graphene $1|-1|1$ system. Our main motivation for studying this particular setup is to obtain knowledge about the ratio of lattice-scale, ultra-short-range anisotropic couplings $u_{z}$ and $u_{xy}$. Such couplings are unable to pick out unique bulk ground states at $\nu=\pm 1$ in the absence of spin and valley Zeeman couplings. Hence we assumed in this work that $E_{Z},E_{V}>0$, which is realistic for experimentally relevant situations. In physical graphene samples it is believed that $u_{z}>0$ while $u_{xy}<0$, and that the two have roughly the same magnitude. Furthermore, their ratio can be altered by Landau-level mixing [65, 66, 67, 68] and the screening environment [69]. Pinning down this ratio and its evolution with screening and other tuning parameters would be invaluable in determining the phases of the $\nu=0$ system, which is yet to be fully understood.
In contrast to $\nu=0$, where the nature of the phase is not fully understood, the ground states at $\nu=\pm 1$ in the physical range of couplings with both spin and valley Zeeman couplings nonzero are known to be valley-polarized spin ferromagnets [45, 51]. This makes it easier to determine the values of the couplings themselves. $E_{Z}$ is known from the total $B$ field, and in principle, one can determine $E_{V}$ by measuring the gap at charge neutrality at $B_{\perp}=0$. A great advantage of this setup is that the in situ tuning parameter $B_{\perp}$ allows us to alter the ratio of $E_{V}$ with respect to the other couplings.
@@
The second important fact in understanding our results is that in the physical region of parameters, $u_{z}>0,~{}u_{xy}<0$, when $u_{z}>|u_{xy}|$ the system prefers to break the spin $U(1)$ symmetry at the interfaces, while in the opposite case $u_{z}<|u_{xy}|$ the system prefers to break the valley $U(1)$ symmetry (for some particular ${B_{\perp}}$). Since the interfaces between $\nu=1$ and $\nu=-1$ can very roughly be thought of as miniature regions of $\nu=0$, this is consistent with the fact that in the corresponding regions the $\nu=0$ bulk ground state breaks exactly those symmetries.
@@
Keeping these two facts in mind, we can easily understand the cases that we considered in Section V with $u_{z}>|u_{xy}|$. There is no valley rotation in these cases, and thus the magnon transmission drops to zero at high energy. For instance, Fig. 5 shows the HF interface structure for the case $u_{z}^{0}=4,~{}u_{xy}^{0}=-3$ and $E_{V}=5$. One can see that there is no valley rotation for any value of $B_{\perp}$. The corresponding magnon transmission is shown in Fig. 6, with the transmission dropping to zero at high energies. Similarly for $E_{V}=1$ the HF structure of the interfaces is shown in Fig. 10, with no valley rotation for any $B_{\perp}$. The corresponding magnon transmission is shown in Fig. 11, and shows vanishing transmission at high energies. The more interesting case is $u_{z}<|u_{xy}|$, believed to occur for unscreened or lightly screened samples [69]. Here a crucial role is played by the valley Zeeman field $E_{V}$. For vanishing $E_{V}\to 0$, the physics of magnon transmission was analyzed by Wei et al[47], and it was found that there is nearly perfect transmission at high energy. Focusing on $E_{V}$ moderate to large, at small $B_{\perp}$, $E_{V}>|u_{\alpha}|,E_{z}$, the configurations at the interfaces do not break any symmetries, and the magnon transmission vanishes at high energies. This is illustrated in panels (a) and (b) of Fig. 7. The corresponding magnon transmission vs energy is shown in panels (a) and (b) of Fig. 8. However, there is a threshold $B_{\perp}$ at which the couplings $u_{\alpha}$ become dominant over $E_{V}$, beyond which the interfaces break the valley $U(1)$ symmetry (Fig. 7(c)), restoring nearly perfect transmission at high energies as shown in Fig. 8(c). This threshold $B_{\perp}$ depends not only on the couplings, but also on the width of the intermediate $\nu=-1$ region, and can be estimated in our model given the sample geometry. This is one of the main results of our work, because this threshold field provides quantitative information about the coupling constants.
In this context, the earlier experiments discussed in the introduction on $1|-1|1$, are not exactly in the regime we study, and so it would be very helpful to have detailed experimental studies at varying $B_{\perp}$ and varying sizes of the intermediate region for the $1|-1|1$ system. So far, the most recent investigation has been restricted to fillings $10<\nu<$ 1 for the intermediate region [29], which, while physically very interesting, is much more complicated than the $1|-1|1$ system that we analyze.
@@
Let us turn to some of the assumptions that underlie our approach. We have assumed ultra-short-range interactions throughout. Consequently, to choose unique bulk ground states at $\nu=\pm 1$, we assumed nonzero spin and valley Zeeman couplings. Relaxing the ultra-short-range assumption in the physical region of parameters $u_{z}>0,~{}u_{xy}<0$ does not change the phases of $\nu=\pm 1$, as has been shown in previous work [45, 51]. However, introducing interactions beyond ultra-short-range does produce new phases at $\nu=0$. Because of our sharp interfaces, the $\nu=0$ regions here are fairly narrow, and we believe interactions beyond USR will not have any qualitative effect on our results. Secondly, we have focused on incident magnons with $q_{y}=0$. In Appendix A, we show some results for incident magnons with $q_{y}\neq 0$, which look qualitatively similar to our results in Section V. Thirdly, we have ignored disorder and finite temperature effects. Disorder can induce the magnons to scatter elastically, thereby reducing the transmission. At $T\neq 0$, thermally generated collective modes will be present in the system, and could scatter the electrically generated magnons inelastically.
@@
There are many open questions that could in principle be addressed by a detailed analysis of magnon transmission. The most important is the $\nu=0$ state, which is yet to be completely understood. It is believed that Landau level mixing leads to the interactions acquiring a range of the magnetic length $\ell$ [41, 42, 69]. Introducing such interactions leads to the appearance of new phases which manifest the coexistence of CAF and bond order, and are separated from the bond-ordered and CAF states by second-order phase transitions [41, 42, 44]. It should be possible to vary the screening to make the system traverse this coexistence phase. Presumably, the magnon transmission properties of this phase differ from that of the standard CAF phase. Furthermore, fractional quantum Hall phases near $\nu=0$ also display a rich variety of phases in the physical regime of parameters, which could be explored via magnon transmission [29]. Another interesting future direction could be exploring shot noise related measurements [70, 71, 72, 73, 74] for the magnon transmission, which could help to reveal the topological orders of the underlying fractional quantum hall systems. We hope to address these and other such questions in the near future.
@@ Acknowledgements.
The authors would like to thank Nemin Wei and Chunli Huang for illuminating discussions. SJD would like to acknowledge both Harish-Chandra Research Institute, Prayagraj, and the cluster computing facilities at the International Center for Theoretical Sciences (ICTS), Bangalore. He also acknowledges ICTS for hospitality and visitor funding. SR and GM would like to thank the VAJRA scheme of SERB, India for its support under grant number VJR/2017/000114. GM thanks the Aspen Center for Physics (NSF Grant No. PHY-2210452 and PHY-160761) for its hospitality.